OpenStudy (anonymous):

Can I get help with my Cal 1 quiz

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok i will post the questions gives me a sec

OpenStudy (amistre64):

the term "quiz" tends to be a touchy subject.

OpenStudy (amistre64):

but other than that, this does look like a doozey

OpenStudy (anonymous):

tell me about it.

OpenStudy (amistre64):

i think that top thing is a u sub

OpenStudy (anonymous):

yes...

OpenStudy (amistre64):

the x on the outside is a derivative of the inside

OpenStudy (amistre64):

with some modifications that is

OpenStudy (amistre64):

u = 4+2x^n du/dx = 2n x^(n-1) du/2n = x^(n-1) dx

OpenStudy (amistre64):

where n is considered to be constant

OpenStudy (anonymous1):

For the second integral you could use u-substitution with u = x²+7x-4.

OpenStudy (anonymous1):

Then du/dx = 2x+7.

OpenStudy (amistre64):

doh!! whyd i take the hard one lol

OpenStudy (anonymous):

How can i set this up can someone draw it out fully fo me

OpenStudy (amistre64):

\[\frac{\int u^{1/2}du}{2n}\] then integrate like normal

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

the final results you really need to obtain on your own tho

OpenStudy (amistre64):

gives you that sense of accomlishment :) and its in a normal looking integration that you should be able to manipulate now

OpenStudy (anonymous):

ok cool thanks

OpenStudy (amistre64):

and .... if you get something other than a for the first one, try again lol